How many grams of silver sulfate can be dissolved in 450 ml of 0.200M sodium sulfate soultion? Ksp Ag2SO4 1.2x10^-5
Answer #1

This is the common ion effect.

Ag2SO4------>2Ag+ + SO42-
     x                   2x         x
Ksp = [Ag+]^2[SO4^]
substitute the given values

1.2*10-5 = (2x) ^2*x = 4x3   

X = 1.732*10-2

Let the solubility of silver sulfate, Ag2SO4, in 0.2 M sodium sulfate, Na2SO4, be X M

Ksp = [Na+]^2[SO4^]

1.2*10-5 = (2x) ^2*(x+0.2)

X = 1.1419

Hence the solubility of silver sulfate, Ag2SO4, in 0.2 M sodium sulfate, Na2SO4 is 1.1419/311.8 = 0.0036 g/L

Therefore for 450 ml, solubility of Ag2SO4 = 0.0036/450 = 8.1*10-6 g